# Training Set Reconstruction from Differentially Private Forests: How Effective is DP?

### v2

@@ 4 Experimental Study
We now empirically evaluate our proposed reconstruction attack, and investigate whether it succeeds in accurately inferring a RF training data despite the DP protection. The source code and data needed to reproduce all our experiments and figures are available at https://github.com/vidalt/DRAFT-DP, under a MIT license.
4.1 Experimental Setup
@@
Datasets and target models. We rely on three datasets, using the same data preprocessing as in Ferry et al. [2024b]. The COMPAS dataset [Angwin et al., 2016] contains records of $7{,}206$ criminal offenders from Florida, each characterized by $15$ binary attributes, and is used for recidivism prediction. The UCI Adult Income dataset [Dua and Graff, 2017] gathers data on $48{,}842$ individuals from the 1994 U.S. census. Each individual is described by $20$ binary features, with the objective of predicting whether an individual earns more than \$50K/year. Finally, the Default of Credit Card Clients (Default Credit) dataset [Yeh and hui Lien, 2009] includes $29{,}986$ customer records from Taiwan, with $22$ binary attributes, to predict payment defaults.
@@
The $\varepsilon$-DP RFs are trained as described in Section 2.2. In practice, we build upon the efficient and modular implementation of IBM’s diffprivlib library [Holohan et al., 2019], which we slightly adapt for our experiments. Following Ferry et al. [2024b], each experiment uses a randomly sampled training set of $N=100$ examples. While this sample size is in line with the literature and allows all experiments to be conducted in a moderate time frame, we also report results with larger training datasets in Appendix E, demonstrating the scalability of the approach. We build DP RFs of size $\lvert\mathcal{T}\rvert\in\{1,5,10,20,30\}$ and of maximum depth $d\in\{3,5,7\}$. Empirical evidence from Fan et al. [2003] suggests that the predictive accuracy of a random forest typically plateaus beyond $30$ trees, with $10$ to $30$ trees being sufficient for robust performance. This is particularly true in the DP setting, in which adding trees consumes additional privacy budget [Fletcher and Islam, 2019]. Finally, we evaluate privacy parameters $\varepsilon\in\{0.1,1,5,10,20,30\}$, ranging from very tight to large privacy budgets, which reflect commonly used values in practice [Desfontaines, 2021].
Reconstruction attack. We use the OR-Tools CP-SAT solver [Perron and Didier, 2024] (v9) to solve the CP formulation introduced in Section 3. Each model resolution is limited to a maximum of $2$ hours of CPU time using $16$ threads with up to $7$ GB of RAM per thread. Note that in practice, feasible solutions – i.e., reconstructions that are compatible with the RF’s structure and the observed noisy counts – can be obtained using only a small fraction of the $2$-hours timeout (in average, less than two minutes), as reported in Appendix C.1. All experiments are run on a computing cluster over a set of homogeneous nodes with AMD EPYC 7532 (Zen 2) @ 2.40 GHz CPU.
We consider two baselines in our experiments. As in Ferry et al. [2024b], a random reconstruction baseline is used to quantify the amount of information that can be inferred without knowledge of the RF (i.e., only based on the number of examples $N$, the different attributes $M$, their domains and one-hot encoding). This baseline randomly guesses the value of each attribute of each example while remaining consistent with one-hot encoding when applicable (we average its performance over $100$ runs). To quantify the effect of DP on the reconstruction error, we also run the state-of-the-art DRAFT attack against the trained RFs without the DP protection (i.e., with the true leaves’ counts).
In line with the literature [Ferry et al., 2024b], while a reconstructed dataset is inferred, first it is aligned with the original training set using a minimum cost matching with the Manhattan distance, in order to identify which reconstructed examples correspond to which original ones. Then, the proportion of binary attributes that differ is averaged over the entire dataset to quantify the reconstruction error. Each experiment is repeated using five different random seeds and the results are averaged.
4.2 Results
Result 1. Reconstruction attacks succeed against DP RFs, even with meaningful privacy guarantees. Table 1 reports the reconstruction error for all our experiments. For commonly protective privacy budgets (e.g., $\varepsilon=1$ or $5$), our attack achieves significantly lower reconstruction error compared to the random baseline. This indicates that the adversary was able to gain insights from the RF to better infer the training data. Nonetheless, the reconstruction error consistently increases when reducing the privacy budget $\varepsilon$, which suggests that DP partially mitigates the reconstruction success. To assess this, Figure 3 compares the effectiveness of our reconstruction to that of DRAFT [Ferry et al., 2024b] applied on the same RF (without DP). For very tight privacy budgets (e.g., small $\varepsilon$), DP significantly increases the reconstruction error compared to the DRAFT baseline. However, as $\varepsilon$ increases, the protective effect of DP diminishes, and the difference in reconstruction error between our attack and DRAFT becomes much smaller, even for $\varepsilon$ values commonly considered protective.
Figure 2: Average reconstruction error as a function of the privacy budget $\varepsilon$ used to fit the target RF, for different numbers of depth-5 trees $|\mathcal{T}\rvert$ on the COMPAS dataset. For comparison, we also report the reconstruction error of DRAFT applied to the same RFs without DP protection.
Result 2. DP leads to a complex trade-off between the RFs’ size and the reconstruction attack success. As can be seen in Figure 3, larger DP RFs do not necessarily provide more information to the attacker, unlike non-DP RFs. A threshold effect is observed: the reconstruction error decreases as the number of trees increases, up to a point (typically around 5-10 trees), before increasing again. This behavior is consistent across all privacy budgets and is likely due to the privacy budget being distributed across more trees, resulting in noisier counts for individual leaves. Similarly, while deeper trees in the non-DP setting provide more information to the attacker, this is not always the case for DP RFs (see also, Table 1) since smaller leaves’ counts are proportionally more affected by the noise.
Result 3. Protection against reconstruction attacks comes at the expense of predictive accuracy. As shown in Table 1, the considered DP mechanism effectively mitigates the success of reconstruction attacks at very low $\varepsilon$ values (e.g., $\varepsilon=0.1$). However, at these privacy levels, the resulting RFs have poor predictive accuracy, often performing no better than a majority prediction classifier, as illustrated in Figure 3. Notably, only RFs with trivial predictive performance provide no advantage to the reconstruction adversary (compared to the random reconstruction baseline). This is also visible in Figure 4, in which most RFs either perform better than a majority classifier but provide a significant advantage to the reconstruction attacker (upper left corner), or provide no advantage to the attacker but perform worse than a majority classifier (lower right corner).
@@ 4.2 Results
Table 1: Average reconstruction error for different numbers of trees $\lvert\mathcal{T}\rvert$ , tree depths $d$ and privacy budgets $\varepsilon$ for the three datasets. Reconstruction errors worse than the random baseline are bold. Cases where the solver failed to find a feasible reconstruction within the time limit are marked with “–”.
@@
$\lvert\mathcal{T}\rvert=10$ $\varepsilon=0.1$ 0.28 0.31 0.34 0.25 0.28 0.28 0.32 0.31 0.30
$\varepsilon=1$ 0.17 0.19 0.23 0.12 0.14 0.16 0.21 0.24 0.24
$\varepsilon=5$ 0.15 0.12 0.15 0.07 0.07 0.06 0.15 0.16 0.19
$\varepsilon=10$ 0.13 0.10 0.11 0.06 0.05 0.04 0.14 0.13 0.13
@@
$\lvert\mathcal{T}\rvert=30$ $\varepsilon=0.1$ 0.34 0.35 0.33 0.33 0.36 0.38

### v4

@@ IV Experimental Study
We now empirically evaluate our proposed reconstruction attack, and investigate whether it succeeds in accurately inferring a RF training data despite the DP protection. The source code and data needed to reproduce all our experiments and figures are available at https://github.com/vidalt/DRAFT-DP.
IV-A Experimental Setup
@@
Datasets. We rely on three datasets, using the same data preprocessing as in [9]. The COMPAS dataset [28] contains records of 7,2067{,}206 criminal offenders from Florida, each characterized by $15$ binary attributes, and is used for recidivism prediction. The UCI Adult Income dataset [29] gathers data on 48,84248{,}842 individuals from the 1994 U.S. census. Each individual is described by $20$ binary features, with the objective of predicting whether an individual earns more than \$50K/year. Finally, the Default of Credit Card Clients dataset [30] (Default Credit) contains 29,98629{,}986 customer records from Taiwan, each with $22$ binary attributes, and is used to predict payment defaults.
@@
Target models. The DP RFs are trained as described in Section II-B. In particular, we note that they satisfy pure $\varepsilon$-DP, which is therefore the focus of our experiments. In practice, we build upon the efficient and modular implementation of IBM’s diffprivlib library [21], which we slightly adapt for our experiments. Following [9], each experiment uses a randomly sampled training set of $N=100$ examples. While this sample size is in line with the literature and allows all experiments to be conducted in a moderate time frame, we also report results with larger training datasets in Appendix E, demonstrating the scalability of the approach. We build DP RFs of size $\lvert\mathcal{T}\rvert\in\{1,5,10,20,30\}$ and of maximum depth $d\in\{3,5,7\}$. Empirical evidence from [31] suggests that the predictive accuracy of a random forest typically plateaus beyond $30$ trees, with $10$ to $30$ trees being sufficient for robust performance. This is particularly true in the DP setting, in which adding trees consumes additional privacy budget [15]. Finally, we evaluate privacy parameters $\varepsilon\in\{0.1,1,5,10,20,30\}$, ranging from very tight to large privacy budgets, which reflect commonly used values in practice [25].
@@
TABLE I: Average reconstruction error for different numbers of trees $\lvert\mathcal{T}\rvert$ , tree depths $d$ and privacy budgets $\varepsilon$ for the three datasets. Reconstruction errors worse than the random baseline are bold. Cases where the solver failed to find a feasible reconstruction within the time limit are marked with “–”.
@@
$\lvert\mathcal{T}\rvert=10$ $\varepsilon=0.1$ 0.28 0.29 0.34 0.25 0.27 0.28 0.32 0.31 0.30
$\varepsilon=1$ 0.17 0.20 0.23 0.12 0.13 0.16 0.21 0.24 0.24
$\varepsilon=5$ 0.15 0.12 0.15 0.07 0.06 0.06 0.15 0.16 0.19
$\varepsilon=10$ 0.13 0.10 0.11 0.06 0.05 0.04 0.14 0.14 0.13
@@
$\lvert\mathcal{T}\rvert=30$ $\varepsilon=0.1$ 0.34 0.35 - 0.33 0.33 - 0.36 0.38 -
@@
In Appendix C-A, we compare the predictive performance of the considered DP RFs (for a large privacy budget $\varepsilon=30$) with that of standard, non-DP RFs built using the popular CART implementation from the scikit-learn [23] Python library. This comparison aims to separate the effect of the DP RF algorithmic design from the impact of DP itself. These design choices include omitting bootstrap sampling and selecting split values and attributes at random rather than using an information gain criterion. Our results show that, with identical hyperparameters, DP RFs with $\varepsilon=30$ achieve test accuracies comparable to those of traditional RFs, thereby confirming the soundness of the algorithmic design.
Reconstruction attack. We use the OR-Tools CP-SAT solver [27] (v9) to solve the CP formulation introduced in Section III. Each model resolution is limited to a maximum of 22 hours of CPU time using $16$ threads with up to $7$ GB of RAM per thread. Note that in practice, feasible solutions (i.e., reconstructions that are compatible with the RF’s structure and the observed noisy counts) can be obtained using only a small fraction of the $2$-hours timeout (in average, less than two minutes), as reported in Table VI in the Appendix C-C. All experiments are run on a computing cluster over a set of homogeneous nodes with AMD EPYC 7532 (Zen 2) @ 2.40 GHz CPU.
Baselines. We consider two baselines in our experiments. As in [9], a random reconstruction baseline is used to quantify the amount of information that can be inferred without knowledge of the RF (i.e., only based on the number of examples $N$, the different attributes $M$, their domains and one-hot encoding). This baseline randomly guesses the value of each attribute of each example while remaining consistent with one-hot encoding when applicable (we average its performance over $100$ runs). Note that since the reconstruction error is computed over the encoded attributes (provided as inputs to the RF), consistency with respect to one-hot-encoding naturally leads to reconstruction errors lower than $0.5$, even for the random baseline. To quantify the effect of DP on the reconstruction error, we also run the state-of-the-art DRAFT attack against the trained RFs without the DP protection (i.e., with the true leaves’ counts).
Reconstruction error measure. In line with the literature [9], while a reconstructed dataset is inferred, first it is aligned with the original training set using a minimum cost matching with the Manhattan distance, in order to identify which reconstructed examples correspond to which original ones. To do so, the Manhattan distance between each reconstructed and original example is computed. The resulting distance matrix then instantiates a minimum weight matching in bipartite graphs, also known as linear sum assignment problem, which we solve using the Scipy [32] Python library. Then, the proportion of binary attributes that differ is averaged over the entire dataset to quantify the reconstruction error.
Each experiment is repeated using five different random seeds and the results are averaged.
IV-B Results
We now highlight our key empirical findings and illustrate each of them with a subset of representative results.
Result 1. Reconstruction attacks succeed against DP RFs, even with meaningful privacy guarantees. Table I reports the reconstruction error for all our experiments.
Figure 2: Average reconstruction error as a function of the privacy budget $\varepsilon$ used to fit the target DP RF, for different numbers of depth-5 trees $|\mathcal{T}\rvert$ on the COMPAS dataset. For comparison, we also report the reconstruction error of DRAFT applied to the same RFs without DP protection (the x-axis does not apply for this baseline), using a dashed line.
For commonly protective privacy budgets (e.g., $\varepsilon=1$ or $5$), our attack achieves significantly lower reconstruction error compared to the random baseline. This indicates that the adversary was able to gain insights from the RF to better infer the training data. Nonetheless, the reconstruction error consistently increases when reducing the privacy budget $\varepsilon$, which suggests that DP partially mitigates the reconstruction success. To assess this, Figure 2 compares the effectiveness of our reconstruction to that of DRAFT [9] applied on the same RF (without DP [^3] Note that the original DRAFT attack formulation cannot be directly applied to DP RFs, since any inconsistency between leaf counts (likely to arise from the added random noise) would render the underlying optimization problem infeasible.). For very tight privacy budgets (e.g., small $\varepsilon$), DP significantly increases the reconstruction error compared to the DRAFT baseline. However, as $\varepsilon$ increases, the protective effect of DP diminishes, and the difference in reconstruction error between our attack and DRAFT becomes much smaller, even for $\varepsilon$ values commonly considered protective. Overall, this comparison confirms that the proposed attack asymptotically approaches the original DRAFT method for large privacy budgets $\varepsilon$.
We nevertheless observe from Table I that no feasible reconstructions were found within the given time limit for the largest and deepest ($|\mathcal{T}\rvert=30$ and $d=7$) DP RFs with the tightest privacy budget $\varepsilon=0.1$. We emphasize that this does not indicate that the reconstruction problem is infeasible, as by design, the true DP RF training set is always part of the CP formulation search space with very high probability. Rather, it suggests that exploring the attack search space and identifying feasible reconstructions becomes computationally more challenging and would require an increase of the solver’s running time. Two main factors are at stake. First, increasing either the number of trees or their depth directly increases the number of variables. Second, tighter privacy budgets lead to larger domains for the noise variables. Together, these effects substantially enlarge the search space.
Result 2. DP leads to a complex trade-off between the RFs’ size and the reconstruction attack success. As can be seen in Figure 2, larger DP RFs do not necessarily provide more information to the attacker, unlike non-DP RFs. A threshold effect is observed: the reconstruction error decreases as the number of trees increases, up to a point (typically around 5-10 trees), before increasing again. This behaviour is consistent across all privacy budgets and is likely due to the privacy budget being distributed across more trees, resulting in noisier counts for individual leaves. Similarly, while deeper trees in the non-DP setting provide more information to the attacker, this is not always the case for DP RFs (see also, Table I) since smaller leaves’ counts are proportionally more affected by the noise. These observations highlight the importance of carefully tuning the hyperparameters of DP RFs, as different trade-offs between predictive accuracy and empirical vulnerability to our reconstruction attack can arise even under a fixed privacy budget $\varepsilon$. For example, on the COMPAS dataset and for $\varepsilon=1$, a DP RF composed of $\lvert\mathcal{T}\rvert=10$ depth-33 trees and another composed of $\lvert\mathcal{T}\rvert=5$ depth-77 trees yield approximately the same reconstruction error under our attack (0.12, see Table I). However, their training accuracies differ substantially ($0.57$ versus $0.61$), as reported in Figures 8(a) and 8(c) in Appendix C-D.
Result 3. Protection against reconstruction attacks comes at the expense of predictive accuracy. As shown in Table I, the considered DP mechanism effectively mitigates the success of reconstruction attacks at very low $\varepsilon$ values (e.g., $\varepsilon=0.1$). However, at these privacy levels, the resulting RFs have poor predictive accuracy, often performing no better than a majority prediction classifier (constantly predicting the majority class), as illustrated in Figure 3. Notably, only DP RFs with trivial predictive performance provide no advantage to the reconstruction adversary (compared to the random reconstruction baseline). This is also visible in Figure 4, in which most DP RFs either perform better than a majority classifier but provide a significant advantage to the reconstruction attacker (upper left corner), or provide no advantage to the attacker but perform worse than a majority classifier (lower right corner). This is observed across all considered datasets and maximum depths, as evidenced in the detailed results reported in Appendix C-D. Notably, the only DP FRs not falling within this dichotomy consist of a single tree, which prevents information from being combined across learners. While we included these special cases for completeness, the probabilistic reconstruction attack of [10] would be more appropriate in such settings, although it is limited to single decision trees and does not natively handle DP.